The siren on an emergency vehicle has a frequency of



840 Hz. The emergency vehicle passes you at a speed of



28 m/s .The frequency of the receding siren detected by you is?

1
Expert's answer
2022-01-06T09:07:28-0500

The frequency of the receding siren detected by you is


"f=\dfrac{c}{c+v}f_0=\dfrac{340}{340+28}840=776\text{ Hz}."